# Gibbs states and Brownian models for coexisting haze and cloud droplets

### v1

@@
Abstract
Clouds microphysics describes the formation and evolution of cloud droplets, rain, and ice particles. It is among the most critical factors in determining the cloud’s size, lifetime, precipitation, and radiative effect. Among all cloud types, the small clouds, characterized by weak updrafts, that are close to the haze-to-cloud transition pose challenges in measuring them and understanding their properties. They are superabundant but hard to capture by satellites and often falsely regarded as aerosols. Köhler’s theory explains droplet activation and their growth by condensation at the earliest stages of cloud development. It fully describes the thermodynamic state of a single drop but falls short when explaining the collective behavior of large populations of particles. This is especially important when the supersaturation pool is limited. We present an analytical framework to extend Köhler’s theory to coexisting cloud droplets. Our results suggest hysteresis and asymmetry in the process of droplet activation and deactivation. The turbulent nature of clouds is incorporated into our model formulation as Brownian noise to provide explicit droplet size distributions and activation timescales. The theoretical findings are confronted with experimental data stemming from laboratory clouds created in the Pi convection chamber, suggesting a new way of understanding haze-to-cloud transitions and small cloud formation processes.
@@ INTRODUCTION
Clouds play a key role in the climate energy balances and in the water cycle Bengtsson (2010). A large part of the uncertainties in climate predictions is attributed to limitations in our understanding of cloud physics and therefore, how to parameterize such clouds in global climate models Zelinka et al. (2020). Out of all cloud types, a great part of the cloud research challenges is attributed to small, warm clouds that are highly abounded but hard to measure. Many of these clouds are too small relative to earth-observing satellites’ resolution, and/or their optical signature is too weak Koren et al. (2007); Várnai and Marshak (2009); Eytan et al. (2020).
@@
The cloud’s droplet size distribution (DSD) is among the most important microphysical properties. The DSD properties are coupled to all dynamical, turbulent, optical, and stochastic cloud processes. DSDs affect the cloud’s evolution in time, rain processes, cloud depth, size, and lifetime Khain and Pinsky (2018). When supersaturation is high enough, cloud condensation nuclei (CCNs) are activated into cloud droplets and grow by condensation. This is the main mechanism for droplet growth at the early stages of a warm cloud. If the cloud’s dynamics is weak (shallow clouds) or if the aerosol concentration is high, growth droplets by condensation could be the main mechanism throughout the whole cloud’s lifetime as droplets do not become large and varied enough to support efficient autoconversion of droplets by the stochastic processes of collision and coalescence (Rogers and Yau, 1989, §7-8).
In warm clouds that are controlled by weak updrafts and low supersaturations, only a small part of the aerosol (the large and most hygroscopic) are activated. A competition between the droplet curvature (Kelvin) effect, which suppresses activation, and the solute (Raoult) effect, which reduces the critical supersaturation for activation, Köhler (1936) allows for suspended particles in air parcels to coexist in two main thermodynamical states known as: haze and activated droplet Khain and Pinsky (2018). While hazy particles are in stable equilibrium, activated particles harvest the available water vapor and grow until equilibrium is restored.
@@
The mechanisms of activation and deactivation are, however, not symmetric. In order to activate a cloud droplet, it is necessary to attain a critical level of supersaturation $\lambda_{K}$, typically well above $0\%$. When supersaturation is decreased to $\lambda_{K}$ in the deactivation process, an activated cloud droplet is big enough to still experience diffusion of water vapor onto its surface and, hence, it is expected that this positive feedback makes the threshold for deactivation strictly lower than that of $\lambda_{K}$. Consequently, activation-deactivation undergoes hysteresis, which has been previously reported in the cloud physics literature Korolev and Mazin (2003); Arabas and Shima (2017). From a theoretical perspective, in the work of Arabas and Shima (2017) it was shown on analytical grounds that the activation process of a cloud particle undergoes a saddle-node bifurcation and, when coupled to supersaturation evolution, hysteresis is numerically observed in a parcel model and attributed to kinetic limitations and the presence of a cusp bifurcation (Kuznetsov, 2004, §8.2).
Another level of complexity in the determination of microphysical properties concerns clouds living in eminently turbulent environments. Indeed, turbulence heavily influences the life cycle of clouds; see Shaw (2003) for a review on the topic. In particular, the incorporation of turbulent models provides the means to explain droplet size spectra broadening Sardina et al. (2015); Grabowski and Abade (2017), as well as describe the mixing of the cumulus clouds with their exterior Abade, Grabowski, and Pawlowska (2018); Eytan et al. (2020) and parameterize mixed-phase clouds in global circulation models Field et al. (2014); Furtado et al. (2016). In the context of warm clouds, localized temperature gradients and turbulence at the smallest scales provoke supersaturation fluctuations that allow activation of haze particles even in mean-subsaturated environments Prabhakaran et al. (2020), and similarly, cloud particles susceptible of deactivating Abade, Grabowski, and Pawlowska (2018). Thus, microphysical fluctuations are a key feature in obtaining faithful descriptions of droplet size spectra.
@@
This research explores the activation and deactivation of cloud droplets in warm clouds, offering new theoretical perspectives. We aim to find analytical formulas for describing the size distribution of droplets (DSDs) when both small cloud droplets and haze particles coexist, assuming condensation as the only growth mechanism. Similar to previous work Arabas and Shima (2017), we analyze how droplet growth due to condensation changes dramatically as the surrounding air’s supersaturation level varies. To the Köhler instability we add a sink term in growth rate which aims at (i) modeling the supersaturation budget, and (ii) parameterising other stagnating mechanisms into the condensational growth, for instance, due to sedimentation Krueger (2020) or the presence of a thermal inversion layer that reduces supersaturation and stops droplet growth Hirsch et al. (2014); Altaratz et al. (2021). The resulting equation describing condensation growth for droplets of a uniform size allows for multiple stable states to co-exist besides the ”haze” state, meaning the system exhibits multistability. Although the haze state remains stable, this multistability suggests that activated droplets can get stuck around a specific size due to the balancing act between two factors: the Köhler instability and mechanisms that decrease supersaturation or remove particles; see also Korolev and Mazin (2003); Arabas and Shima (2017); Chandrakar et al. (2020).
@@
However, turbulence effects throw a wrench in the picture. The stable states become “fuzzy states” that are encountered statistically (metastable) at various rates of occurrence. This means that particles can switch between activated and deactivated states, similar to how Brownian particles move around in a bumpy landscape. These turbulent jolts also significantly impact the activation-deactivation process into the form of various possible hysteresis loops (see Fig. 5 below). As a result, droplets can become activated at lower levels of supersaturation compared to the standard Köhler threshold.
@@
Finally, our framework allows us to derive analytical formulas for the size distribution of droplets (DSDs) within a cloud volume, accounting for these stochastic turbulent effects. In particular, our stochastic model predicts that droplet size distributions in Pi cloud convection chamber experiments behave like Gibbs states. It is shown that this prediction closely matches the actual DSDs observed in these experiments (see Fig. 6 below).
@@ RESULTS
The theory of Köhler is the thermodynamical backbone to explain the formation of cloud droplets out of suspended aerosols in the atmosphere. It establishes the critical relative humidity threshold beyond which suspended particles activate and grow by diffusion (condensation) Köhler (1936). The critical threshold is above saturation (hence, the supersaturation term is used), and it depends on the initial size and chemical composition of the particle in question. These particles are, according to Köhler, present in two distinct/dichotomical thermodynamic states: The first is known as haze and it is observed when the aerosol is humidified, although remains at equilibrium with its moist environment. The second state is referred to as activated droplet, whereby the particle harvests the available moisture by diffusion and, in principle, grows indefinitely. However, in nature, the available water molecules that are in supersaturation are limited, and in many cases, droplets and haze coexist in the cloud, and particles can oscillate between the two states. Particularly, in cases of small, warm clouds Hirsch et al. (2017); Altaratz et al. (2021).
The critical relative humidity threshold needed for activation of a given aerosol with a solubility constant $k$ and a dry radius $r_{d}$ is analytically obtained using the Köhler’s equation Köhler (1936); Rogers and Yau (1989). Here, to simplify some derivations, $\lambda$ denotes the part of the relative humidity above $100\%$. If $r(t)$ denotes the radius of a given droplet at time $t$, its evolution is approximated by Köhler (1936); Arabas and Shima (2017):
@@
where $D$ is the diffusivity constant, $A$ relates to the water surface tension, $B=kr_{d}^{3}$, $\lambda$ is the ambient supersaturation and $\lambda_{eq}(r)$ determines the equilibrium supersaturation at the surface of a droplet of radius $r$; commonly known as Köhler curve Wallace and Hobbs (2006). Equation (1) shows that the particle is in equilibrium when $\lambda=\lambda_{eq}(r)$. To simplify the proposed model, we replace $r$ with the variable $X=r^{2}/2D$, which has time-units and could be viewed as relaxation time for diffusion. With this change of variable, Eq. (1) translates into the nondimensional form:
@@
with $f(X)$ denoting the equilibrium supersaturation $\lambda_{eq}(r)$ as function of $X$. A rudimentary analysis of $f(X)$ reveals a single global maximum located at $X_{K}=3B/(2DA)$, and with critical supersaturation value $\lambda_{K}=f(X_{K})=(4A^{3}/27B)^{1/2}$, see also (Wallace and Hobbs, 2006, §6.1). An example of Köhler curve is shown in Fig. 1a) whose parameters are given in Table 1. This curve reaches its maximum at $X_{K}=6\cdot 10^{-3}\mathrm{s}$, or at $r^{2}=0.48\mathrm{\mu m}^{2}$.
Table 1: Köhler curve’s parameters for NaCl.
@@
For a given dry radius $r_{d}$, when the supersaturation $\lambda$ exceeds the corresponding $\lambda_{K}$, the particle is activated into a cloud droplet and grows by diffusion. On the contrary, for a given ambient supersaturation $\lambda$, particles whose radii are less than $(2DX_{u})^{1/2}$ remain stable in a haze state; see Fig. 1a).
While Köhler theory is established for single particles, clouds consist of large families of particles coexisting in the same humidity environment. When a collection of droplets consumes the available supersaturation, the particles may cease growing by condensation, and the droplet size distributions (DSDs) stagnate and stop being displaced towards larger sizes. Here we modify the Köhler equation to the case of many particles in conditions of small warm clouds, when the available supersaturation is a limit factor. The proposed model extends the condensational growth equation—governed by the Köhler curve— by coupling all particles to supersaturation consumption and by adding stochastic disturbances, to account for (micro) small-scale turbulent effects.
Figure 1: Panel a): Köhler curve associated with $\mathrm{NaCl}$; see Table 1 for the parameter values. A dry radius of $50\mathrm{nm}$ is prescribed. $\lambda_{K}$ denotes Köhler’s critical supersaturation and $\lambda$, an example of supersaturation. Panel b): In black, the Köhler curve associated with $\mathrm{NaCl}$ as Panel a). The cyan to blue curves show different M-Köhler curves associated with Eq. (6) for different values the parameter $\beta(N)$ as indicated in the legend. For the cyan curve are shown furthermore its intersection with a given supersaturation $\lambda$ in between $\lambda_{c}$ and $\lambda_{h}$. Each filled dot corresponds to a stable equilibrium while the the empty dot corresponds to the unstable equilibrium. Panel c): The potential function associated with the cyan M-Köhler curve for the supersaturation value as indicated in the title. The solid/empty dots locate the stable/unstable equilibria.
@@
In cases with relatively small supersaturation, as particles grow, they may consume the available supersaturation and therefore reduce their growth rate until reaching stagnation. In the realm of single-particle models these effects can be modeled by the introduction of a sink function into Eq. (1):
@@
in which $g$ denotes the sink term. The function $f-g$ is called a M-Köhler curve hereafter. The term $g$ should remain negative and should decrease as $X$ increases. The sink term in growth rate aims at (i) modeling the supersaturation budget, and (ii) parameterizing other stagnating mechanisms into the condensational growth, for instance, due to sedimentation Krueger (2020) or the presence of a thermal inversion layer that reduces supersaturation and stops droplet growth Hirsch et al. (2014); Altaratz et al. (2021).
The sink term can also relate to first-order fluctuations in supersaturation. Indeed, when a sizable amount of particles coexist in the same cloud parcel, the constant supersaturation assumption has to be amended taking account changes in $\lambda$ in the course of time, assuming $\lambda$ in Eq. (3) to be time-dependent. At a first order, we are concerned with describing the evolution of $\lambda(t)=\lambda+\lambda^{\prime}(t)$ with $\lambda^{\prime}(0)=0$, i.e. we are seeking for a model of the rate of change of the fluctuations $\lambda^{\prime}$ around a given value $\lambda$.
Following Arabas and Shima (2017), the rate of change $\lambda^{\prime}$ can be approximated by the ratio $-\dot{\rho_{v}}/\rho_{vs}$, in which $\rho_{v}$ denotes the ambient vapour density away from the droplet while, $\rho_{vs}$, denotes the vapour density at saturation. Then, by expressing $\rho_{v}$ as $\rho_{v}=4\pi N\rho_{w}r^{3}/3$, accounting for the density of liquid water $\rho_{w}$ and the droplet volume concentration $N$ of identical (monodisperse) particles in the cloud parcel, we arrive at:
$\dot{\lambda^{\prime}}=-\beta(N)\dot{\left(X^{\frac{3}{2}}\right)},$ (4)
@@
see (Arabas and Shima, 2017, Eqns. (9) and (10)).
The integration of Eq. (4) yields to $\lambda(t)=\lambda-\beta(N)X^{3/2}$ which eventually leads to the following condensational growth model:
$\dot{X}=\lambda-f(X)-\beta(N)X^{3/2}.$ (6)
@@
Sufficient conditions for Eq. (7) to have two positive solutions are easily derivable. Actually, a simple proposition inferring the existence of two positive steady states $X^{\ast}_{h}$ and $X^{\ast}_{c}$ (with $X^{\ast}_{c}\geq X^{\ast}_{h}$) to condensational growth model analogues to Eq. (6) for a sink function $-\beta(N)X^{\alpha}$ ($\alpha>1$) is derived in in Methods (see Proposition 1 therein). In any event, as shown in Fig. 1b) the locations of $X^{\ast}_{h}$ and $X^{\ast}_{c}$ depend on the strength of the sink function’s coefficient $\beta(N)$.
We define then the following supersaturation values $\lambda_{h}=f(X^{\ast}_{h})-g(X^{\ast}_{h})$, which corresponds to the local maximum of $f-g$, and $\lambda_{c}=f(X^{\ast}_{c})-g(X^{\ast}_{c})$, corresponding to its local minimum. In particular, $\lambda_{h}\geq\lambda_{c}$.
Due to the presence of this new local minimum (compared to the Köhler curve), as supersaturation $\lambda$ varies, the number of equilibria varies through two saddle-node bifurcations, as sumarized in Fig. 2.
Figure 2: Going through multiple equilibria as $\lambda$ varies. Two saddle-node (SN) bifurcations occur, one at $\lambda=\lambda_{c}$ and one $\lambda=\lambda_{h}$. For $\lambda<\lambda_{c}$, there is only a hazy stable equilibrium $X_{h}$ (blue dot in the leftmost upper diagram). As one crosses from below the SN bifurcation point at $\lambda=\lambda_{c}$, two other equilibria emerge for a total of 3 equilibria coexisting for $\lambda_{c}<\lambda<\lambda_{h}$ (middle upper diagram): an activated stable equilibrium $X_{c}$ (red dot), an unstable equilibrium $X_{u}$ (open circle), and a haze equilibrium (blue dot). At the SN bifurcation occurring at $\lambda=\lambda_{h}$, $X_{u}$ and $X_{h}$ collide (purple dot on the rightmost lower diagram), and only the activated equilibrium survives as $\lambda>\lambda_{h}$ (red dot in the rightmost upper diagram).
These two saddle-node bifurcations provide the nonlinear backbone for hysteresis to take place as discussed in Hysteresis effects, below. These hysteresis effects are well-known to be revealed in presence of stochastic disturbances; see e.g. Berglund (2011). We discuss below the physical motivation of introducing such disturbances in Eq. (6) or more generally, Eq. (3).

### v2

@@
Abstract.
Cloud microphysics studies include how tiny cloud droplets grow, and become rain. This is crucial for understanding cloud properties like size, lifespan, and impact on climate through radiative effects. Small, weak-updraft clouds near the haze-to-cloud transition are especially difficult to measure and understand. They are abundant but hard to capture by satellites. Köhler’s theory explains initial droplet growth but struggles with large particle groups. Here, we present a stochastic, analytical framework building on Köhler’s theory to account for (monodisperse) aerosols and cloud droplets interaction through competitive growth in a limited water vapor field. These interactions are modeled by sink terms while fluctuations in supersaturation affecting droplet growth are modeled by nonlinear, white noise terms. Our results identify hysteresis mechanisms in the droplet activation and deactivation processes. Our approach allows for multimodal cloud’s droplet size distributions supported by lab experiments, offering a new perspective on haze-to-cloud transition and small cloud formation.
Keywords. Köhler Theory |||| Droplet Size Distributions |||| Stochastic Modeling |||| Hysteresis |||| Gibbs States
@@ Introduction
Clouds play a key role in the climate energy balances and in the water cycle [Ben10]. A large part of the uncertainties in climate predictions is attributed to limitations in our understanding of cloud physics and therefore, how to parameterize such clouds in global climate models [ZMM+20]. Out of all cloud types, small, warm clouds, are highly abundant and pose serious challenges for cloud research. Their diminutive size often falls below the resolution limits of Earth-observing satellites, making direct observation difficult. Additionally, their weak optical signatures can hinder detection and measurement [KRK+07, VM09, EKA+20]. These factors contribute to a substantial knowledge gap regarding the role and impact of small, warm clouds in the global climate system.
@@
The cloud’s droplet size distribution (DSD) is among the most important microphysical properties. The DSD properties are coupled to all dynamical, turbulent, optical, and stochastic cloud processes. DSDs affect the cloud’s evolution in time, rain processes, cloud depth, size, and lifetime [KP18]. When supersaturation is high enough, cloud condensation nuclei (CCNs) are activated into cloud droplets and grow by condensation. This is the main mechanism for droplet growth at the early stages of a warm cloud. If the cloud’s dynamics is weak (shallow clouds) or if the aerosol concentration is high, condensation could be the main growth mechanism throughout the whole cloud’s lifetime as droplets do not become large and varied enough to support efficient autoconversion of droplets by the stochastic processes of collision and coalescence; see Chapters 7 and 8 in [RY89].
Köhler theory provides a framework for understanding the activation of aerosol particles into cloud droplets. In low supersaturation conditions, only a small subset of the aerosol population, typically the largest and most hygroscopic particles, are activated. The activation process is governed by a delicate balance between the Kelvin effect, which opposes activation due to the curvature of the droplet surface, and the Raoult effect, which favors activation due to the presence of dissolved solutes. This interplay is central to Köhler theory [Köh36]. As a result of this competition, aerosol particles can exist in two distinct thermodynamic states: haze particles and activated droplets [KP18]. Haze particles remain typically in stable equilibrium with the surrounding environment, while activated droplets grow by diffusion. However other microphysics regimes in a turbulent flow can lead to intricate interactions between haze and activated droplets, resulting e.g. into nonlinear behaviors such as oscillatory pulses of activation at low particle concentrations [YSS+24].
@@
The mechanisms of activation and deactivation are, however, not symmetric. In order to activate a cloud droplet, it is necessary to attain a critical level of supersaturation $\lambda_{K}$, typically well above $0\%$. When the supersaturation is decreased to $\lambda_{K}$ during the deactivation process, an activated cloud droplet may remain sufficiently large to continue experiencing diffusion of water vapor onto its surface. This positive feedback mechanism results into a deactivation threshold that is strictly lower than $\lambda_{K}$. In this case, activation-deactivation exhibits hysteresis, a phenomenon previously documented in the cloud physics literature [KM03, AS17].
From a nonlinear dynamics perspective, such an hysteresis phenomenon results from the presence of two saddle-node bifurcation curves that meet tangentially (Chapter 8.2 in [Kuz04]). Recall that a saddle-node bifurcation is a type of bifurcation where two equilibrium points “collide” and disappear as a system’s parameter is varied (here supersaturation). It is at the core of the notion tipping points [AWVC12] and occurs in many fields of physics such as combustion theory [BE13, Che18], energy balance climate models [Ghi76, NCCJ81], or ocean dynamics [HSA+11, WCD+19, CLM23, LC23]. In the context of cloud physics, the authors in [AS17] have shown that the activation process of a cloud particle following the Köhler condensational growth equation, undergoes such a saddle-node bifurcation in which the haze equilibrium merges with an unstable equilibrium. At this bifurcation point, the stability of the haze equilibrium is lost due to this merging. On analytical and numerical grounds it was shown in [AS17] that hysteresis does occur for air-parcel models accounting for temperature and pressure variations, suggesting the presence of the other saddle-node bifurcation where the activated equilibrium merges with an unstable one. In this work, we provide a general framework for hysteresis to take place from the Köhler condensational growth equation alone by incorporating a sink term accounting e.g. for the evolution of ambient heat and moisture content, or sedimentation effects.
Our approach also accounts for turbulent effects. Clouds in turbulent environments exhibit a complex interplay between microphysical processes and turbulent dynamics. Turbulence strongly influences the life cycle of clouds, as demonstrated in [Sha03]. Turbulent models are essential for understanding droplet size spectra broadening [SPBC15, GA17], the mixing of cumulus clouds with their surroundings [AGP18, EKA+20], and the parameterization of mixed-phase clouds in global circulation models [FHFK14, FFB+16]. In warm clouds, turbulence-induced localized temperature gradients and small-scale fluctuations can lead to supersaturation variations. These variations can activate haze particles even in environments with a mean subsaturation, and similarly, cloud particles may deactivate in saturated conditions if the fluctuations are sufficiently strong [PSK+20, AGP18]. Therefore, microphysical fluctuations are crucial for accurately representing droplet size spectra in turbulent cloud environments. We represent these effects by means of stochastic parameterizations involving Brownian motions [CSG11, LC23] and possibly depending on the particle’s size.
@@
Thus, this research explores the activation and deactivation of cloud droplets in warm clouds, offering new theoretical perspectives. We aim to find analytical formulas for describing the cloud’s DSDs when both small cloud droplets and haze particles coexist, assuming condensation as the only growth mechanism. To this end, we add a sink term to the classical formulation of Köhler’s equation in order to parameterize the interaction of a population of droplets or other size-depleting mechanisms, like water vapor consumption or sedimentation [AS17, HKAA15, Kru20]. The resulting equation describing condensational growth of droplets composing a monodisperse aerosol population allows for multiple stable states to co-exist besides the “haze" state, meaning the system exhibits multistability. Although the haze state remains stable, this multistability suggests that activated droplets can get stuck around a specific size due to the balancing act between two factors: the Köhler instability and mechanisms that decrease supersaturation or remove particles; see also [KM03, AS17, CSY+20].
@@
This is where turbulence effects throw a wrench in the picture. The stable states become “fuzzy states” that are encountered statistically (metastable) at various rates of occurrence. This means that particles can switch between activated and deactivated states, similar to how Brownian particles move around in a bumpy landscape. These turbulent jolts also strongly impact the activation-deactivation process into the form of various possible hysteresis loops; see Section "Hysteresis effects" below. As a result, droplets can become activated at lower levels of supersaturation compared to the standard Köhler threshold.
@@
Finally, our framework allows us to derive analytical formulas for the DSDs within a cloud volume, accounting for these stochastic turbulent effects. In particular, our stochastic model predicts that the DSDs in the Pi cloud convection chamber experiments behave like Gibbs states, i.e. an equilibrium probability distribution that maximizes the system’s entropy subject to constraints on its average energy like in statistical mechanics [JKO98]. It is shown that our theoretical predictions based on this formalism closely match the actual DSDs observed in these experiments; see Section "Pi-Chamber empirical distributions vs Gibbs states" below.
@@ Results
The theory of Köhler is the thermodynamical backbone to explain the formation of cloud droplets out of suspended aerosols in the atmosphere. It establishes the critical relative humidity threshold beyond which suspended particles activate and grow by diffusion (condensation) [Köh36]. The critical threshold is above saturation (hence, the supersaturation term is used), and it depends on the initial size and chemical composition of the particle in question. These particles are, according to Köhler, present in two distinct/dichotomical thermodynamic states: The first is known as haze and it is observed when the aerosol is humidified, although it remains at equilibrium with its moist environment. The second state is referred to as activated droplet, whereby the particle harvests the available moisture by diffusion and, in principle, grows indefinitely. However, in nature, the available water vapor molecules that are in supersaturation are limited, and in many cases, droplets and haze coexist in the cloud, and particles can oscillate between the two states. Particularly, in cases of small, warm clouds [HKA+17, AKA+21].
The critical relative humidity threshold needed for activation of a given aerosol with a solubility constant $k$ and a dry radius $r_{d}$ is analytically obtained using Köhler’s equation [Köh36, RY89]. Here, to simplify some derivations, $\lambda$ denotes the part of the relative humidity above $100\%$. If $r(t)$ denotes the radius of a given droplet at time $t$, its evolution is approximated by [Köh36, AS17]:
@@
where $D$ is the diffusivity constant, $A$ relates to the water surface tension, $B=kr_{d}^{3}$, $\lambda$ is the ambient supersaturation and $\lambda_{eq}(r)$ determines the equilibrium supersaturation at the surface of a droplet of radius $r$; commonly known as Köhler curve [WH06]. Equation (1) shows that the particle is in equilibrium when $\lambda=\lambda_{eq}(r)$. To simplify the proposed model, we replace $r$ with the variable $X=r^{2}/2D$, which has time-units and could be viewed as relaxation time for diffusion. With this change of variable, Eq. (1) translates into the nondimensional form:
@@
with $f(X)$ denoting the equilibrium supersaturation $\lambda_{eq}(r)$ as function of (the scaled squared radius) $X$. A rudimentary analysis of $f(X)$ reveals a single global maximum located at $X_{K}=3B/(2DA)$, and with critical supersaturation value $\lambda_{K}=f(X_{K})=(4A^{3}/27B)^{1/2}$, see also Chapter 6.1 in [WH06]. An example of Köhler curve is shown in Fig. 1a) whose parameters are given in Table 1. This curve reaches its maximum at $X_{K}=6\cdot 10^{-3}\mathrm{s}$, or at $r^{2}=0.48\mathrm{\mu m}^{2}$.
Table 1. Köhler curve’s parameters for NaCl.
@@
For a specified dry radius, $r_{d}$, a particle will activate into a cloud droplet when the ambient supersaturation, $\lambda$, surpasses its critical supersaturation, $\lambda_{K}$. This activation process occurs through diffusional growth. Conversely, for a given supersaturation less than $\lambda_{K}$, particles with radii smaller than $(2DX_{u})^{1/2}$ will remain in a stable haze state. These equilibrium points are visually represented by the intersection of a horizontal line at the level of $\lambda$ and the Köhler curve, as depicted in Figure 1a. The solid and empty dots within this figure denote the stable and unstable equilibria, respectively.
While Köhler theory is established for single particles, clouds consist of large families of particles coexisting in the same humidity environment. When a collection of droplets consumes the available supersaturation, the particles may cease growing by condensation, and the DSDs stagnate and stop being displaced towards larger sizes [KM03, SBK17]. Here we modify the Köhler equation to the case of many particles in conditions of small warm clouds, when the available supersaturation is a limit factor. The proposed model extends the condensational growth equation—governed by the Köhler curve— by coupling all particles to supersaturation consumption and by adding stochastic disturbances, to account for (micro) small-scale turbulent effects.
Figure 1. Köhler and M-Köhler curves, and underlying potential. Panel a): Köhler curve associated with $\mathrm{NaCl}$; see Table 1 for the parameter values. A dry radius of $50\mathrm{nm}$ is prescribed. $\lambda_{K}$ denotes Köhler’s critical supersaturation and $\lambda$, an example of supersaturation. Panel b): In black, the Köhler curve associated with $\mathrm{NaCl}$ as Panel a). The cyan to blue curves show different M-Köhler curves associated with Eq. (6) for different values the parameter $\beta(N)$ as indicated in the legend. For the cyan curve are shown furthermore its intersection with a given supersaturation $\lambda$ in between $\lambda_{c}$ and $\lambda_{h}$. Each filled dot corresponds to a stable equilibrium while the empty dot corresponds to the unstable equilibrium. Panel c): The potential function associated with the cyan M-Köhler curve for the supersaturation value as indicated in the title. The solid/empty dots locate the stable/unstable equilibria.
@@
In cases with relatively small supersaturation, as particles grow, they may consume the available supersaturation and therefore reduce their growth rate until reaching stagnation. In the realm of single-particle models, these effects can be modeled by the introduction of a sink function into Eq. (2):
@@
in which $g$ denotes the sink term. The function $f-g$ is called a M-Köhler curve hereafter. The term $g$ should remain negative and should decrease as $X$ increases, as its main role is for accounting for saturation mechanisms in the particle growth. More precisely, this sink term in growth rate aims to (i) model the supersaturation budget, and (ii) parameterize other stagnating mechanisms into the condensational growth rate, for instance, due to sedimentation [Kru20] or the presence of a thermal inversion layer that reduces supersaturation and stops droplet growth [HKL+14, AKA+21].
The sink term can also relate to first-order fluctuations in supersaturation due to condensation. Indeed, when a sizable amount of particles coexist in the same cloud parcel, the constant supersaturation assumption has to be amended taking account changes in $\lambda$ in the course of time, assuming $\lambda$ to be time-dependent in the first equation of Eq. (2). At a first order, we are concerned with describing the evolution of $\lambda(t)=\lambda+\lambda^{\prime}(t)$ with $\lambda^{\prime}(0)=0$, i.e. we are seeking for a model of the rate of change of the fluctuations $\lambda^{\prime}$ around a given value $\lambda$.
Following [AS17], the rate of change $\lambda^{\prime}$ can be approximated by the ratio $-\dot{\rho_{v}}/\rho_{vs}$, in which $\rho_{v}$ denotes the ambient vapor density away from the droplet while, $\rho_{vs}$, denotes the vapor density at saturation. Then, by expressing $\rho_{v}$ as $\rho_{v}=4\pi N\rho_{w}r^{3}/3$, accounting for the density of liquid water $\rho_{w}$ and the droplet number concentration $N$ of identical (monodisperse) particles in the cloud parcel, we arrive at:
$\frac{\,\mathrm{d}}{\,\mathrm{d}t}\lambda^{\prime}=-\beta(N)\frac{\,\mathrm{d}}{\,\mathrm{d}t}{\left(X^{\frac{3}{2}}\right)},$ (4)
@@
see Eqns. (9) and (10) in [AS17].
By integrating each side of Eq. (4), we obtain $\lambda^{\prime}(t)=-\beta(N)X^{3/2}$. When this expression of $\lambda^{\prime}$ is used in $\lambda(t)=\lambda+\lambda^{\prime}(t)$ we arrive at $\lambda(t)=\lambda-\beta(N)X^{3/2}$ which leads, from the first equation in Eq. (2), to the following condensational growth model:
$\dot{X}=\lambda(t)-f(X)=\lambda-f(X)-\beta(N)X^{3/2}.$ (6)
@@
Sufficient conditions for Eq. (7) to have two positive solutions are easily derivable. Actually, the existence of haze and activated steady states, $X^{\ast}_{h}$ and $X^{\ast}_{c}$ (with $X^{\ast}_{c}\geq X^{\ast}_{h}$), can be inferred for condensational growth models analogue to Eq. (6) with more general sink functions $-\beta(N)X^{\alpha}$ ($\alpha>1$); see Proposition 1 in Section "Multistability" in Material and Methods. In any event, as shown in Fig. 1b) the locations of $X^{\ast}_{h}$ and $X^{\ast}_{c}$ depend on the strength of the sink function’s coefficient $\beta(N)$.
We define then the following supersaturation values $\lambda_{h}=f(X^{\ast}_{h})-g(X^{\ast}_{h})$, corresponding to the local maximum of $f-g$, and $\lambda_{c}=f(X^{\ast}_{c})-g(X^{\ast}_{c})$, corresponding to its local minimum. In particular, $\lambda_{h}\geq\lambda_{c}$.
Due to the introduction of a new local minimum compared to the original Köhler curve (Fig. 1a), the number of equilibria varies with changes in supersaturation, $\lambda$, through two saddle-node bifurcations. As depicted in Figure 2, these bifurcations occur when the haze equilibrium (solid blue dot) approaches and merges with the unstable node (empty circle). Conversely, as $\lambda$ decreases, a similar bifurcation takes place when the activated equilibrium approaches the unstable node. The solid and purple dots in the bottom panels denote the critical values of $\lambda$ at which these bifurcations occur.
Figure 2. Going through multiple equilibria as $\lambda$ varies. Two saddle-node (SN) bifurcations occur, one at $\lambda=\lambda_{c}$ and one $\lambda=\lambda_{h}$. For $\lambda<\lambda_{c}$, there is only a haze stable equilibrium $X_{h}$ (blue dot in the leftmost upper diagram). As one crosses from below the SN bifurcation point at $\lambda=\lambda_{c}$, two other equilibria emerge for a total of 3 equilibria coexisting for $\lambda_{c}<\lambda<\lambda_{h}$ (middle upper diagram): an activated stable equilibrium $X_{c}$ (red dot), an unstable equilibrium $X_{u}$ (open circle), and a haze equilibrium (blue dot). At the SN bifurcation occurring at $\lambda=\lambda_{h}$, $X_{u}$ and $X_{h}$ “collide” (purple dot on the rightmost lower diagram), and only the activated equilibrium survives as $\lambda>\lambda_{h}$ (red dot in the rightmost upper diagram).
These two saddle-node bifurcations serve as the nonlinear foundation for the hysteresis phenomenon discussed in the following Section on "Hysteresis effects." These hysteresis effects are well-known to be revealed in presence of stochastic disturbances combined with slow parameter variations; see e.g. [Ber11]. We discuss below the physical motivation of introducing such disturbances in Eq. (6) or more generally, in Eq. (3).